OpenStudy (anonymous):

What is the value of x? 12 - x = 7 x=_________

OpenStudy (anonymous):

@notcomplying

OpenStudy (anonymous):

x = 19

OpenStudy (anonymous):

Thanks!

Directrix (directrix):

12 - x = 7 -12 -12 ---------------- Add -12 to both sides

OpenStudy (anonymous):

what are u talking about directrix

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

its 19 hes confused

OpenStudy (ipwnbunnies):

It's definitely not 19.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

Wait whatt?

OpenStudy (anonymous):

12 + 7 =19

OpenStudy (anonymous):

its a subtraction sign

OpenStudy (anonymous):

you do the opposite to both sides

OpenStudy (anonymous):

Oh ok

Directrix (directrix):

12 - x = 7 -12 -12 ---------------- -x = -5 x = 5

Directrix (directrix):

@Kyliejones

OpenStudy (anonymous):

Dude wtf shutup

OpenStudy (anonymous):

Ok well lets go to the next question

OpenStudy (anonymous):

What is the value of x? -5x = -35 x=_____

OpenStudy (anonymous):

x = 7

OpenStudy (anonymous):

Alrighty!

OpenStudy (anonymous):

What is the value of x? -2x + 8 + -16